The demand of a good is D=100-9.4P; the supply of it is S=9.2P-35. At the current market price, there is a surplus of a 9.0 units on the market. Please find the current market price.
1
Expert's answer
2020-09-02T12:11:04-0400

If there is a surplus on the market then S>D, and the surplus is calculated, as supply subtracting demand: S-D= 9.

9.2P-35-100+9.4P=9

18.6P=144

P=7.74 - is the current market price